OpenStudy (anonymous):

if p = 2(l + w) then l= ??

OpenStudy (lalaly):

\[l=\frac{p}{2}-w\]

OpenStudy (anonymous):

Divide both sides of the equation by 2 to start. p/2 = l + w Then subtract w from both sides of the equation to get p/2 - w = l or l= p/2 - w